OpenStudy (mtalhahassan2):

Need help!! Solve a) 4^x+6(4^-x)=5

OpenStudy (mtalhahassan2):



OpenStudy (mtalhahassan2):

@Nnesha

OpenStudy (mtalhahassan2):

@mathmate @sorryimfrankie @MTALHAHASSAN2 @randomecorner @preteenpony @MTALHAHASSAN2 @BlossomCa

OpenStudy (randomecorner):

pemdas parenthise exponets multiply divide add subtract

OpenStudy (freckles):

multiply both sides by 4^x

OpenStudy (freckles):

you will have a quadratic in terms of 4^x afterwards

OpenStudy (freckles):

\[4^x+6(4^{-x})=5 \\ \text{ multiply both sides by} 4^{x} \\ 4^{2x}+6(1)=5(4^x) \\ 4^{2x}-5(4^x)+6=0 \\ (4^x)^2-5(4^x)+6=0\]

OpenStudy (freckles):

can you finish from here?